OpenStudy (anonymous):

what is the integral 2 to 5 of abs(x-3)dx . I know you have to split them up but not so sure how and what that does to the problem. also this problem... integral -1 to 2 of (x-2 absx)dx Thanks!!

ganeshie8 (ganeshie8):

yes we need to split the integral such that the function never becomes negative : \[ \large |x-3| = \left\{ \begin{array}{lr} 3-x & : x < 3\\ x-3 & : x \ge 3 \end{array} \right. \]

ganeshie8 (ganeshie8):

\[\large \int \limits_2^5 |x-3| dx = \int \limits_2^3( 3-x) dx + \int \limits_3^5 (x-3) dx \]

ganeshie8 (ganeshie8):

see if that makes more or less sense...

ganeshie8 (ganeshie8):

we have used the definition of |x| function : \[ \large |x| = \left\{ \begin{array}{lr} x & : x < 0\\ -x & : x \ge 0 \end{array} \right. \]

OpenStudy (anonymous):

for the first one, where did you get integral 2 to 3 and 3 to 5 from? and for the second one.. what would the integral look like? \[\int\limits_{-1}^{0} (x-2x) + \int\limits_{0}^{2} (x-2x)dx\] ?

OpenStudy (aravindg):

For first one he just split the integral where there is change in how function is defined. x-3=0 means x=3 So 2 to 5 interval was split as sum of two integrals 2 to 3 then 3 to 5

OpenStudy (agreene):

The second one kinda depends on what it is, your notation is a bit ambiguous, is it: \[\int_{-1}^2(x-2)|x|dx \space \space \text{or} \space \space \space \int_{-1}^2 x-2|x| dx\]

OpenStudy (anonymous):

sorry, it is \[\int\limits_{1}^{2} (x-2\left| x \right|)dx \]

OpenStudy (agreene):

in that case, you can do a lot of splitting: \[\int_{-1}^2x-2|x|dx =\int_{-1}^2xdx-\int_{-1}^22|x|dx=\int_{-1}^2xdx-2\int_{-1}^2|x|dx\] you will still have to fix the continuity issue for x=-1